OpenStudy (anonymous):

What are the solutions of the system? y=x+5 y=x^2-x+2

OpenStudy (anonymous):

either x=-1 y=4 or x=3 y=8

OpenStudy (anonymous):

Thank you so much! can you help me with one more?

OpenStudy (anonymous):

y=x^2-2x-8 y=x+2